# softMathematical software

### v1

@@ 6. Application to entropic region delimitation
    We also asked ourselves whether some of the remaining 154 types of extreme polymatroids can be excluded from being entropic using stronger probabilistic rules from [31, § V.B]. This is indeed the case: we found further 11 types which cannot be probabilistic because their induced CI models ${\mathcal{M}}$ do not satisfy the (extended version of the) probabilistic rule (I:2) saying that
@@
    But only 5 of these 11 types can reliably be excluded from being almost entropic as they do not satisfy the above implication with $M=\emptyset$; these finer conclusions follow from the (method of) rule derivation in [31, § V.B].
@@ 7. Conclusions
{open}
Find all coatoms and irreducible elements of $(\mathfrak{sg}(N),\subseteq)$ in case $|N|=5$.
@@
We also verified computationally that $\mathfrak{sg}^{\diamond}(N)=\mathfrak{sg}^{\diamond\diamond}(N)$ in case $|N|=4$. Computing this “second-order self-adhesion” required a significant implementation effort and took much longer than any of our other computations. Nonetheless, we expect that $\mathfrak{sg}^{\diamond}(N)=\mathfrak{sg}^{\diamond\diamond}(N)$ does not hold for $|N|=5$.
@@
{open}
Find out whether $\mathfrak{sg}^{\diamond}$ admits a finite “axiomatic” characterization, too.
@@ Acknowledgements
Most of this work was inspired by the research of our late colleague Fero Matúš. The authors discussed in person the topic of this paper at two meetings in 2022, one of them was the DAGSTUHL seminar 22301 “Algorithmic Aspects of Information Theory”.
@@ Appendix A Algorithm to compute the canonical basis
This is a pseudo-code of an algorithm to get the canonical generator ${\mathbb{G}}_{{\mathcal{F}}}$ on basis of a Moore family ${\mathcal{F}}$ of subsets of $X$ or its closure operation $\mbox{\rm cl}_{{\mathcal{F}}}$:
@@
Note that in the step 1, when ${\mathcal{F}}=\mathfrak{f}(N)$ for a suitable CI family $\mathfrak{f}$, one typically has $m=1$. By the construction of ${\mathbb{G}}$, we know that ${\mathbb{G}}\subseteq{{\mathcal{F}}}^{\triangleleft}$ and ${{\mathcal{F}}}\subseteq{\mathbb{G}}^{\triangleright}$ in every phase of the algorithm. In the step 4(b), we are sure that ${\mathbb{G}}$ is the implicational generator of ${\mathcal{F}}$ by the criterion mentioned above Example 2.19.
@@ Appendix B Algorithm to test iterated self-adhesivity
Consider a frame $\mathfrak{f}$ closed under copying, marginalization, intersection, lifting and tight replication, such that $\mathfrak{f}\subseteq\mathfrak{sg}$.
The basis of the algorithm is a subroutine to compute the $\mathfrak{f}^{\diamond}$-closure of a given model ^$\hat{{\mathcal{R}}}\in\mathfrak{f}(M)$, which is a modification of the procedure in Lemma 4.6.
Subroutine $\mathfrak{f}^{\diamond}(\hat{{\mathcal{R}}})$ [the input is ^$\hat{{\mathcal{R}}}\in\mathfrak{f}(M)$]
@@
    put $\mathfrak{f}^{\diamond}(\hat{{\mathcal{R}}}):={\mathcal{R}}$ and exit.
Note that in step 2, by Corollary 4 and Lemma 3.18, we need not test sets $L$ of cardinalities $10,1,|M|-$ 1, and $|M|$. As concerns updating of ${\mathcal{R}}$ in 2(b), by Definition 3.3, $\mathfrak{f}^{\diamond}({\mathcal{R}}|L)\in\mathfrak{f}(M)$.
The main procedure is analogous and differs in small details only. The main difference is that Lemma 3.4 is this time applied to $\mathfrak{f}^{\diamond}$ instead of $\mathfrak{f}$. Note that this modification is possible thanks to Lemma 4.1.
Main procedure $\mathfrak{f}^{\diamond\diamond}(\hat{{\mathcal{M}}})$ [the input is ^$\hat{{\mathcal{M}}}\in\mathfrak{f}^{\diamond}(N)$, where $|N|=4$]
@@
    if $^{\mathcal{M}}\neq\hat{{\mathcal{M}}}$ then ^$\hat{{\mathcal{M}}}\not\in\mathfrak{f}^{\diamond\diamond}(N)$ and its closure is $\mathfrak{f}^{\diamond\diamond}(\hat{{\mathcal{M}}}):={\mathcal{M}}$; exit.
Note that, in step 2, by Lemma 4.1 and Corollary 4, we need not test sets $E$ of cardinalities $0,1$, and $4$. On the other hand, since we are not sure whether $\mathfrak{f}^{\diamond}$ is closed under tight replication (see Task 4) we have to test sets $E$ of cardinality $3$. So, the number of sets $E$ in line 2 is at most 10.
To check double self-adhesivity over $4$ variables one needs to apply the subroutine for $|M|=5,6$, in which case the number of sets $L$ in step 2 of the subroutine is at most 50. Thus, to compute $\mathfrak{f}^{\diamond}({\mathcal{R}}|L)$ one applies the $\mathfrak{f}$-closures over at most 10 variables.
@@ Appendix D Irreducible self-adhesive structural models
The types I.IX. involve all 31 coatoms of the lattice. Their notation follows the classification used in [31, Appendix B]. The actual numbers of involved CI statements are as follows:

### v3

@@ 8. Application to entropic region delimitation
    One of the reviewers raised the question of how many of the remaining 154 types are known to be entropic for sure. The best lower bound we are aware of is provided by linearly representable polymatroids which are known to be entropic. The cone generated by linear $5$-polymatroids was characterized in [11] both in terms of linear inequalities and in terms of its extreme rays: they computed 7943 of such extreme rays falling into 162 permutational types. Of these 162 types exactly 138 coincide with extreme rays of the polymatroidal cone. Thus there are only 16 extreme rays of the $5$-polymatroid cone for which it is unknown whether they are entropic or not. We provide the corresponding coatoms of $\mathfrak{st}^{\diamond}(5)$ on our supplementary materials website.
  3. (3)
    Progress on the remaining 16 types can be made using stronger probabilistic rules from [39, § V.B]. This is indeed the case: we found further 11 types which cannot be probabilistic because their induced CI models ${\mathcal{M}}$ do not satisfy the (extended version of the) probabilistic rule (I:2) saying that
@@
    But only 5 of these 11 types can reliably be excluded from being almost entropic as they do not satisfy the above implication with $M=\emptyset$; these finer conclusions follow from the (method of) rule derivation in [39, § V.B]. We leave this observation for future work.
@@ 9. Conclusions
Open problem 5.
We would like to find all coatoms and irreducible elements of $(\mathfrak{sg}(N),\subseteq)$ in case $|N|=5$.
@@
We also verified computationally that $\mathfrak{sg}^{\diamond}(N)=\mathfrak{sg}^{\diamond\diamond}(N)$ in case $|N|=4$. Computing this “second-order self-adhesion” required a significant implementation effort and took much longer than any of our other computations: while the results on single self-adhesions documented in Appendix C took roughly 18 minutes in total to compute, the double self-adhesion $\mathfrak{sg}^{\diamond\diamond}(4)$ required 6 hours. To explain this sudden increase, recall from Appendix B that computing the closure operator with respect to $\mathfrak{f}^{\diamond\diamond}(N)$ requires the closure for $\mathfrak{f}^{\diamond}(M)$ for $|N|\leq|M|\leq 2|N|-2$, which in turn is computed using the closure operators $\mathfrak{f}(O)$ for $|M|\leq|O|\leq 2|M|-2$. In the case of semi-graphoids over $|N|=4$, this requires semi-graphoid closures over ground sets of size up to $10$. At the maximum $|O|=10$, the corresponding Boolean formula used to compute closures has 184 320 clauses in 11 520 variables. We expect that $\mathfrak{sg}^{\diamond}(N)=\mathfrak{sg}^{\diamond\diamond}(N)$ does not hold for $|N|=5$, but an exhaustive computation is out of our reach.
@@
Open problem 6.
We would like to know whether $\mathfrak{sg}^{\diamond}$ admits a finite “axiomatic” characterization, too.
@@ Acknowledgements
Most of this work was inspired by the research of our late colleague Fero Matúš. Tobias Boege was partially supported by the Academy of Finland grant number 323416 and by the Wallenberg Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. The authors discussed in person the topic of this paper at two meetings in 2022, one of them was the DAGSTUHL seminar 22301 “Algorithmic Aspects of Information Theory”. We also thank both reviewers of DAM journal for the valuable comments which helped to improve the quality of presentation of our results.
@@ Appendix A Algorithm to compute the canonical basis
This is a pseudo-code of the algorithm we used to obtain the canonical generator ${\mathbb{G}}_{{\mathcal{F}}}$ on basis of a Moore family ${\mathcal{F}}$ of subsets of a finite set $X$ or its closure operation $\mbox{\rm cl}_{{\mathcal{F}}}$:
@@
Note that in Step 1, when ${\mathcal{F}}=\mathfrak{f}(N)$ for a suitable CI family $\mathfrak{f}$, one typically has $m=1$. By the construction of ${\mathbb{G}}$, we know that ${\mathbb{G}}\subseteq{{\mathcal{F}}}^{\triangleleft}$ and ${{\mathcal{F}}}\subseteq{\mathbb{G}}^{\triangleright}$ in every phase of the algorithm. In Step 4(b), we are sure that ${\mathbb{G}}$ is the implicational generator of ${\mathcal{F}}$ by the criterion mentioned above Example 2.13.
What is described above is just one of basic versions of the procedure to compute the canonical basis. There are surely various more refined and efficient versions of this procedure [16, § 3.3.2]; the paper [13] was particularly devoted to the comparison of computational complexities of such algorithms.
@@ Appendix B Algorithm to test iterated self-adhesivity
Consider a CI frame $\mathfrak{f}$ closed under copying, marginalization, intersection, lifting, tight replication, and ascetic extension.
The basis of the algorithm is a subroutine to compute the $\mathfrak{f}^{\diamond}$-closure of a given model ^$\widehat{{\mathcal{R}}}\in\mathfrak{f}(M)$, which is a modification of the procedure in Lemma 5.4.
Subroutine $\mathfrak{f}^{\diamond}(\widehat{{\mathcal{R}}})$ [the input is ^$\widehat{{\mathcal{R}}}\in\mathfrak{f}(M)$]
@@
    put $\mathfrak{f}^{\diamond}(\widehat{{\mathcal{R}}}):={\mathcal{R}}$ and exit.
Note that in Step 2, by Corollary 4.12 and Lemma 4.16, we need not test sets $L$ of cardinalities $10,1,|M|-$ 1, and $|M|$. As concerns updating of ${\mathcal{R}}$ in Step 2(b), by Definition 4.3, $\mathfrak{f}^{\diamond}({\mathcal{R}}|L)\in\mathfrak{f}(M)$.
The main procedure is analogous and differs in small details only. The main difference is that Lemma 4.5 is this time applied to $\mathfrak{f}^{\diamond}$ instead of $\mathfrak{f}$. Note that this modification is possible thanks to Lemma 5.1(iii).
Main procedure $\mathfrak{f}^{\diamond\diamond}(\widehat{{\mathcal{M}}})$ [the input is ^$\widehat{{\mathcal{M}}}\in\mathfrak{f}^{\diamond}(N)$, where $|N|=4$]
@@
    if $^{\mathcal{M}}\neq\widehat{{\mathcal{M}}}$ then ^$\widehat{{\mathcal{M}}}\not\in\mathfrak{f}^{\diamond\diamond}(N)$ and its closure is $\mathfrak{f}^{\diamond\diamond}(\widehat{{\mathcal{M}}}):={\mathcal{M}}$; exit.
Note that, in Step 2, by Lemma 5.1(i)(ii) and Corollary 4.12, we need not test sets $E$ of cardinalities $0,1$, and $4$. On the other hand, since we are not sure whether $\mathfrak{f}^{\diamond}$ is closed under tight replication (see Open problem 1) we have to test sets $E$ of cardinality $3$. So, the number of sets $E$ in line 2 is at most 10.
To check double self-adhesivity over $4$ variables one needs to apply the subroutine for $|M|=5,6$, in which case the number of sets $L$ in step 2 of the subroutine is at most 50. Also, to compute $\mathfrak{f}^{\diamond}({\mathcal{R}}|L)$ one applies the $\mathfrak{f}$-closures over at most 10 variables.
@@ Appendix D Irreducible self-adhesive structural models
The types I.IX. involve all 31 coatoms of the lattice. Their notation follows the classification used in [39, Appendix B]. The actual numbers of involved elementary CI statements are as follows: